IDENTIFICATION OF THE [9.1] ^{2}Π_{3/2} - X ^{2}Π_{3/2} TRANSITION OF NICKEL CHLORIDE

\underline{LEAH OBRIEN} and SINAN TUMTURK, Department of Chemistry, Southern Illinois University, Edwardsville, IL 62026-1652.

The near infrared system of NiCl at 9,100 cm^{−1}has been recorded and identified as the [9.1] ^{2}Π_{3/2} - X ^{2}Π_{3/2} transition of NiCl. Molecular constants for the new [9.1] ^{2}Π_{3/2} state will be reported. A comparison of the known energy levels for nickel halides will be presented.